# Retinex-RAWMamba: Bridging Demosaicing and Denoising for Low-Light RAW Image Enhancement

### v1

@@ I $=Convs_{1,5,3}\{cat[\textbf{{X}},sap(\textbf{{X}})]\}$ (8)
The details of DAF are shown in Fig. 2 (b), previous feature map will be firstly concatenated with current feature map at the same level, and this result will be multiplied with previous feature map after the convolution, then it will pass through a convolution with a residual addition. And we can get the fused feature map after a final convolution. Specifically, for the tow feature maps $\textbf{{pre}}\in{\mathbb{R}}^{H\times W\times C}$ and $\textbf{{cur}}\in{\mathbb{R}}^{H\times W\times C}$, they will be fused as following:
@@
Experiments
Datasets and Experiments Environments
SID Dataset
For Sony subset, there are totally 1865 raw image pairs in the training set. Each pair of images contains a short exposure and a long exposure, the short exposure is used as noisy raw, and the long exposure is used as $GT_{raw}$. The original size of all images is $2848\times 4256$. Limited by GPU memory, the data is preprocessed before training, first pack into $4\times 1424\times 2128$, then randomly crop a patch with shape $4\times 512\times 512$ as the input with random data augmentation, such as horizontal/vertical flipping. For the test set, we referred to the DNF(Jin et al. 2023a) settings and deleted the three misaligned scene images.
@@
MCR Dataset
The MCR (Dong et al. 2022) dataset contains 4980 images with a resolution of $1280\times 1024$, including 3984 low-light RAW images, 498 monochrome images (not be used for us) and 498 sRGB images. With indoor and outdoor scenes, different exposure times are set, 1/256s to 3/8s for indoor scenes and 1/4096s to 1/32s for outdoor scenes. And we obtained the raw ground truth as DNF (Jin et al. 2023a) did. The preprocessing is similar to SID dataset, but we don’t randomly crop a patch as the input.
Implementation Details
During training, the batch size is 1 and the initial learning rate is 1e-4, and we use the cosine annealing strategy to reduce it to 1e-5 at the 200th epoch. Adamw optimizer is used and the betas parameter is [0.9,0999] and the momentum is 0.9. The training and testing is completed by a NVIDIA 3090 (24G), A40 (48G),respectively due to the limitation of GPU memory. We also provide the code of merging test on a 24G GPU. Note that the results of merging test are a litter bit smaller than that testing with whole image. And we use PSNR, SSIM (Wang et al. 2004) and LPIPS (Zhang et al. 2018) as the quantitative evaluation metrics.
Comparison with State-of-the-Arts
We conduct experiments on SID (Chen et al. 2018) dataset including Sony and Fuji subsets and MCR (Dong et al. 2022) dataset, and compare with previous SOTA methods including SID (Chen et al. 2018), DID (Maharjan et al. 2019), SGN (Gu et al. 2019), EEMEFN (Zhu et al. 2020), LDC (Xu et al. 2020), LLPackNet (Lamba, Balaji, and Mitra 2020), RRT (Lamba and Mitra 2021), MCR (Dong et al. 2022), RRENet (Huang et al. 2022) and DNF (Jin et al. 2023a).
@@
Table 2: Quantitative results on MCR (Dong et al. 2022) dataset. The top-performing result is highlighted in bold, while the second-best is shown in underline. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better.
@@
Ablation Studies
To demonstrate the validity of our proposed method, we perform ablation experiments on the SID Sony dataset. We first propose a baseline model that consists only of SDB and the unmodified naive visual mamba and GFM from DNF (Jin et al. 2023a). Tab. 3 shows the results of adding or replacing the corresponding module based on the baseline, where RRAWM stands for replacing naive Mamba with RRAWMamba, RDM stands for adding RDM module, and DAF stands for replacing GFM with DAF module. All the ablation experiments were conducted in the same environment.
First, we replaced the baseline’s naive Mamba with the proposed RRAWMamba. The results showed increases of 0.41 dB in PSNR and 0.012 in SSIM, demonstrating that our RRAWMamba, with its eight-directional scanning mechanism, performs well in the demosaicing task. Next, we incorporated the proposed RDM for denoising and automatic exposure correction. The results indicated that although SSIM did not improve, PSNR increased by an additional 0.27 dB. This suggests that the initial exposure of the images was indeed problematic, and our RDM effectively enhances denoising and exposure correction. Finally, we replaced all GFM components in the network with our proposed DAF to improve the stability of the training process. This led to further gains, with PSNR and SSIM increasing by 0.04 dB and 0.001, respectively.
Additionally, we performed a simple visualization of the ablation study, as shown in Fig. 5. The use of RRAWMamba in the baseline enhances dark image details, though brightness remains insufficient. Retinex-RawMamba with RDM addresses this issue, and it is evident that our proposed method achieves better results in both detail and brightness in extremely dark areas.
@@
Baseline RRAWM RDM DAF PSNR SSIM

### v2

@@ 3.6 Domain Adaptive Fusion
The details of DAF are shown in Fig. 2 (c), previous feature map will be firstly concatenated with current feature map at the same level, and this result will be multiplied with previous feature map after the convolution, then it will pass through a convolution with a residual addition. And we can get the fused feature map after a final convolution. Specifically, for the two feature maps $\textbf{{pre}}\in{\mathbb{R}}^{H\times W\times C}$ and $\textbf{{cur}}\in{\mathbb{R}}^{H\times W\times C}$, they will be fused as following:
@@ 3.7 Loss Function
Table 1: Quantitative results of RAW-based LLIE methods on the Sony and Fuji subsets of SID. The top-performing result is highlighted in red, while the second-best is shown in green. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better. ‘-’ indicates the result is not available.
Category Method Venue #Params.(M) Sony Fuji
Single-Stage SID CVPR2018 7.7 28.96 0.787 0.356 26.66 0.709 0.432
DID ICME2019 2.5 29.16 0.785 0.368 - - -
SGN ICCV2019 19.2 29.28 0.790 0.370 27.41 0.720 0.430
LLPackNet BMVC2020 1.2 27.83 0.755 0.541 - - -
RRT CVPR2021 0.8 28.66 0.790 0.397 26.94 0.712 0.446
Multi-Stage EEMEFN AAAI2020 40.7 29.60 0.795 0.350 27.38 0.723 0.414
LDC CVPR2020 8.6 29.56 0.799 0.359 27.18 0.703 0.446
MCR CVPR2022 15.0 29.65 0.797 0.348 - - -
RRENet TIP2022 15.5 29.17 0.792 0.360 27.29 0.720 0.421
DNF CVPR2023 2.8 30.62 0.797 0.343 28.71 0.726 0.391
Ours - 6.2 30.76 0.810 0.328 29.02 0.743 0.382
4 Experiments
4.1 Datasets and Experiments Environments
4.1.1 SID Dataset
For Sony subset, there are totally 1865 raw image pairs in the training set. Each pair of images contains a short exposure and a long exposure, the short exposure is used as noisy raw, and the long exposure is used as $GT_{raw}$. The original size of all images is $2848\times 4256$. Limited by GPU memory, the data is preprocessed before training, first pack into $4\times 1424\times 2128$, then randomly crop a patch with shape $4\times 512\times 512$ as the input with random data augmentation, such as horizontal/vertical flipping. For the test set, we referred to the DNF[17] settings and deleted the three misaligned scene images.
@@ 4.1.1 SID Dataset
4.1.2 MCR Dataset
The MCR [7] dataset contains 4980 images with a resolution of $1280\times 1024$, including 3984 low-light RAW images, 498 monochrome images (not be used for us) and 498 sRGB images. With indoor and outdoor scenes, different exposure times are set, 1/256s to 3/8s for indoor scenes and 1/4096s to 1/32s for outdoor scenes. And we obtained the raw ground truth as DNF [17] did. The preprocessing is similar to SID dataset, but we don’t randomly crop a patch as the input.
4.1.3 Implementation Details
During training, the batch size is 1 and the initial learning rate is 1e-4, and we use the cosine annealing strategy to reduce it to 1e-5 at the 200th epoch. Adamw optimizer is used and the betas parameter is [0.9,0.999] and the momentum is 0.9. The training and testing is completed by a NVIDIA 3090 (24G), A40 (48G),respectively due to the limitation of GPU memory. We also provide the code of merging test on a 24G GPU. Note that the results of merging test are a litter bit smaller than that testing with whole image. And we use PSNR, SSIM [36] and LPIPS [47] as the quantitative evaluation metrics.
4.2 Comparison with State-of-the-Arts
We conduct experiments on SID [5] dataset including Sony and Fuji subsets and MCR [7] dataset, and compare with previous SOTA methods including SID [5], DID [26], SGN [12], EEMEFN [50], LDC [42], LLPackNet [20], RRT [19], MCR [7], RRENet [15] and DNF [17].
@@ 4.2 Comparison with State-of-the-Arts
4.3 Ablation Studies
Table 2: Quantitative results on MCR [7] dataset. The top-performing result is highlighted in red, while the second-best is shown in green. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better.
@@ 4.3 Ablation Studies
To demonstrate the validity of our proposed method, we perform ablation experiments on the SID Sony dataset. We first propose a baseline model that consists only of SDB and the unmodified naive visual mamba in MambaIR [13] and GFM in DNF [17]. Tab. 3 shows the results of adding or replacing the corresponding module based on the baseline, where RAWM stands for replacing naive Mamba with RAWMamba, RDM stands for adding RDM module, and DAF stands for replacing GFM with DAF module. All the ablation experiments were conducted in the same environment.
First, we replaced the baseline’s naive Mamba with the proposed RAWMamba. The results showed increases of 0.41 dB in PSNR and 0.012 in SSIM, demonstrating that our RAWMamba, with its eight-directional scanning mechanism, performs well in the demosaicing task. Next, we incorporated the proposed RDM for denoising and automatic exposure correction. The results indicated that although SSIM did not improve, PSNR increased by an additional 0.27 dB. This suggests that the initial exposure of the images was indeed problematic, and our RDM effectively enhances denoising and exposure correction. Finally, we replaced all GFM components in the network with our proposed DAF to improve the stability of the training process. This led to further gains, with PSNR and SSIM increasing by 0.06 dB and 0.001, respectively.
Furthermore, we conducted a straightforward visualization of the ablation study, depicted in Fig. 5. The incorporation of RAWMamba into the baseline model effectively reduces noise and mitigates the green color distortion. Comparatively, the Retinex-RawMamba approach demonstrates superior color correction capabilities and attains the highest PSNR and SSIM scores. This clearly indicates that our proposed method outperforms others in terms of both detail preservation and color accuracy.
@@
Baseline RAWM RDM DAF PSNR SSIM